# MA-RAG: Multi-Agent Retrieval-Augmented Generation via Collaborative Chain-of-Thought Reasoning

### v1

@@ Abstract
We present MA-RAG, a Multi-Agent framework for Retrieval-Augmented Generation (RAG) that addresses the inherent ambiguities and reasoning challenges in complex information-seeking tasks. Unlike conventional RAG methods that rely on either end-to-end fine-tuning or isolated component enhancements, MA-RAG orchestrates a collaborative set of specialized AI agents: Planner, Step Definer, Extractor, and QA Agents, to tackle each stage of the RAG pipeline with task-aware reasoning. Ambiguities may arise from underspecified queries, sparse or indirect evidence in retrieved documents, or the need to integrate information scattered across multiple sources. MA-RAG mitigates these challenges by decomposing the problem into subtasks, such as query disambiguation, evidence extraction, and answer synthesis, and dispatching them to dedicated agents equipped with chain-of-thought prompting. These agents communicate intermediate reasoning and progressively refine the retrieval and synthesis process. Our design allows fine-grained control over information flow without any model fine-tuning. Crucially, agents are invoked on demand, enabling a dynamic and efficient workflow that avoids unnecessary computation. This modular and reasoning-driven architecture enables MA-RAG to deliver robust, interpretable results. Experiments on multi-hop and ambiguous QA benchmarks demonstrate that MA-RAG outperforms state-of-the-art training-free baselines and rivals fine-tuned systems, validating the effectiveness of collaborative agent-based reasoning in RAG.
@@ 1 Introduction
Recent advances in natural language processing have driven the development of Retrieval-Augmented Generation (RAG) models, which aim to enhance the factual accuracy and contextual relevance of generated text by integrating external knowledge sources (Lewis et al., 2020; Guu et al., 2020; Izacard and Grave, 2021; Lin et al., 2024). These systems address core limitations of Large Language Models (LLMs), such as outdated knowledge (Zhang et al., 2023b; Kasai et al., 2023) and poor generalization to domain-specific queries (Siriwardhana et al., 2023; Xiong et al., 2024), by retrieving top-$k$ documents from an external corpus (Formal et al., 2022; Izacard et al., 2022; Wang et al., 2022a) to ground the model’s output in relevant evidence.
Prior research in RAG has largely concentrated on optimizing three key components—retrieval, augmentation, and generation (Gao et al., 2024; Fan et al., 2024) (Figure 1(a)). Retrieval strategies span sparse methods (Jones, 1972; Robertson and Zaragoza, 2009) and dense retrieval (Reimers and Gurevych, 2019; Karpukhin et al., 2020), each with respective weaknesses such as lexical gaps (Berger et al., 2000) or retrieval failure on out-of-distribution and multi-hop queries (Dai et al., 2023). Augmentation methods often rely on post-retrieval processing such as re-ranking or document summarization (Chen et al., 2020; Glass et al., 2022; Ma et al., 2024) (Figure 1(b)) to improve input quality for the LLM, but add latency and may still fail to filter irrelevant or misleading evidence. More sophisticated approaches introduce iterative retrieval or query rewriting (Jiang et al., 2023b; Asai et al., 2024) (Figure 1(c)), but commonly assume the input query is well-formed and overlook the broader reasoning process across the pipeline. Across retrieval, augmentation, and generation, most existing methods treat components in isolation, failing to resolve ambiguities and reasoning gaps that span multiple stages, such as vague queries, incomplete retrievals, or scattered evidence, ultimately limiting robustness and transparency in complex QA scenarios.
@@ 2 Related Works
Large Language Models (LLMs) has spurred significant advancements over the past few years. Beginning with GPT-1 (Radford et al., 2018) on the Transformer architecture (Vaswani et al., 2017), subsequent models like GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), and the latest GPT-4 (OpenAI, 2024) have significantly enhance capabilities in text understanding and generation. Beyond the GPT series, models such as Mistral (Jiang et al., 2023a), Gemini (Gemini Team, 2023), and LLaMA ((Touvron et al., 2023a), Touvron et al. (2023b)) demonstrate robust performance across various tasks like question-answering and entity recognition (Zhao et al., 2023). Training LLMs involves unsupervised pre-training, supervised fine-tuning, and alignment with human feedback, yet challenges persist in domain-specific tasks (Kandpal et al., 2023). Techniques like PEFT (Houlsby et al., 2019a) optimize fine-tuning efficiency, with emerging methods such as prompt-based learning (Lester et al., 2021; Li and Liang, 2021), adapters (Houlsby et al., 2019b; Fu et al., 2021; Wang et al., 2022b; He et al., 2022), and reparameterization (Hu et al., 2022; Edalati et al., 2022; Dettmers et al., 2023) showing promise by focusing on selective parameter adjustment for enhanced performance.
Retrieval-Augmented Generation (RAG) improves LLM performance by incorporating external knowledge through document retrieval (Lewis et al., 2020; Guu et al., 2020). Despite its promise, RAG faces ongoing challenges in determining what, when, and how to retrieve (Gao et al., 2024). Early efforts integrate retrieval into next-token prediction (Khandelwal et al., 2020; Ram et al., 2023; Liu et al., 2024b) or propose end-to-end pipelines (Guu et al., 2020; Borgeaud et al., 2022; Izacard et al., 2023; Zhang et al., 2024a), while others study knowledge representation and retrieval robustness (Xu et al., 2024; Sarthi et al., 2024). Supervised and contrastive approaches, however, struggle with scalability and domain transfer (Dai et al., 2023; Zhang et al., 2023b; Shi et al., 2024). More structured retrieval approaches include HippoRAG (Gutiérrez et al., 2024), which integrates knowledge graphs for better guidance.
Recent advancements focus on deeper retrieval-reasoning integration. RA-DIT (Lin et al., 2024) retrofits LLMs with retrieval by independently tuning the model for context use and the retriever for relevance. Speculative RAG (Wang et al., 2025) drafts answers using a specialist LM and verifies them with a generalist LM. CD-LM (Li et al., 2025b) accelerates inference via chunk-level retrieval, enabling training-free adaptation. Auto-GDA (Leemann et al., 2025) addresses domain adaptation with synthetic data generation and filtering. Graph-based methods such as ToG-2 (Ma et al., 2025) and SubgraphRAG (Li et al., 2025a) enhance retrieval by leveraging subgraph structures. Reinforcement learning approaches refine retrieval and generation strategies, including episodic memory (Shinn et al., 2023), retrieval policy optimization (Kulkarni et al., 2024), evidence citation (Menick et al., 2022), and reflection-based refinement (Asai et al., 2024; Zhou et al., 2023; Gao et al., 2025), though often at high computational cost. In contrast to these methods and prior agent-based frameworks that rely on heavy supervision, such as RankRAG (Yu et al., 2024), our proposed MA-RAG offers a lightweight, training-free agentic solution.
@@
LLM-based Agentic Systems coordinate multiple specialized agents to solve complex tasks through structured interaction (Guo et al., 2024). Agents operate in varied environments, such sandboxed, physical, or abstract (Hong et al., 2024; Mao et al., 2025; Park et al., 2023), and assume roles that are predefined, emergent, or data-driven (Du et al., 2024; Xiong et al., 2023). Communication follows cooperative, competitive, or debate-based paradigms, structured through centralized or decentralized channels (Liu et al., 2024c; Hong et al., 2024). Agent capabilities are developed through environmental feedback, memory retrieval, or self-evolution (Wang et al., 2023, 2024; do Nascimento et al., 2023; Zhang et al., 2023a; Chen et al., 2024a, b). Recent work has extended agentic designs to RAG. For example, Agentic RAG for time series (Ravuru et al., 2024) routes queries through a hierarchical agent structure, while CollEX (Schneider et al., 2025) enables multimodal retrieval via vision-language agents. In contrast, MA-RAG employs a set of lightweight, specialized agents that collaborate through chain-of-thought reasoning to handle ambiguity, improve transparency, and enhance performance in complex QA settings without any fine-tuning.
@@ 3 Method
3.1 Preliminaries
Retrieval-augmented generation leverages a large corpus of documents or contexts—such as Wikipedia—to provide grounded knowledge for question answering. Given a query $q$ and a corpus $\mathcal{C}$, a dense retriever $\mathcal{R}$ retrieves the top-$k$ relevant contexts $C_{q}=\{c_{1},...,c_{k}\}$. In standard RAG pipelines, a large language model (LLM) generates the answer based on a prompt that includes the query and the retrieved documents:
@@ 3.2 Multi-Agent System for Retrieval-Augmented Generation (MA-RAG)
Planner Agent. The planner initiates the process by analyzing the input query $q$ to perform both query disambiguation and task decomposition. It first identifies potential ambiguities or underspecified elements in the query and reformulates them into clearer sub-questions if necessary. Then, for complex or multi-hop questions, it outputs a structured plan $P=\{s_{1},s_{2},\ldots,s_{n}\}$, where each $s_{i}$ denotes a reasoning subtask. The planner is guided by chain-of-thought prompting with few-shot examples, ensuring interpretable step-wise decomposition that supports grounded and disambiguated reasoning in downstream modules.
@@
Extractor Agent. Retrieved passages often contain redundant or irrelevant content. Instead of appending entire chunks, the extractor performs fine-grained selection and aggregation of sentences or spans that are directly aligned with the current subquery. This process not only filters out noise and mitigates the lost-in-the-middle issue (Liu et al., 2024a), but also enables effective evidence aggregation by combining complementary information from multiple sources into a concise, task-specific evidence set for the QA agent. This aggregated context facilitates more accurate and informed answer generation.
@@ 4.2 Results
Figure LABEL:fig:comparision provides a visual comparison between MA-RAG and several baseline models across multiple datasets. The full numerical results with additional discussion are available in the Appendix. Key observations from our experiments include:
@@
When scaling up to larger models, MA-RAG (Llama3-70B and GPT-4o-mini) outperforms the strongest 70B-scale models, such as ChatQA-1.5 70B and RankRAG 70B, setting new state-of-the-art results on multiple benchmarks. Notably, MA-RAG achieves a score of 59.5 on NQ, 87.2 on TriviaQA, 52.1 on HotpotQA, and 47.5 on 2WikimQA. In particular, on more challenging, multi-hop, and long-tailed datasets like HotpotQA and 2WikimQA, MA-RAG demonstrates significant gains over previous methods. These improvements suggest that the fine-grained query decomposition and passage extraction capabilities inherent in MA-RAG are particularly advantageous in handling complex retrieval conditions. Overall, these results highlight the critical role of multi-agent coordination in improving open-domain QA performance, emphasizing that the integration of specialized agents for different reasoning steps leads to more effective and efficient utilization of external knowledge sources.
@@ 4.3 Ablation Study
Impact of Agents To understand the contribution of each agent in our framework, we conduct an ablation study by removing either the extractor or the planner agent from MA-RAG. Table 1 presents the performance across five QA benchmarks under each setting. When the extractor is removed, the system feeds retrieved documents directly into the prompt without filtering out irrelevant content, resulting in a consistent drop in performance across all datasets. This confirms the critical role of the extractor in refining the input and enhancing the grounding of the model’s responses. On the other hand, removing the planner agent effectively reduces MA-RAG to a single-turn RAG system with document filtering but no query decomposition. While this configuration still performs reasonably well on simpler, single-hop datasets, it fails to handle multi-hop questions that require structured reasoning and intermediate steps. Notably, the most significant performance drop occurs on multi-hop datasets, emphasizing the importance of the planner in guiding the reasoning process for complex queries. These results demonstrate that both agents are essential for robust and adaptable RAG behavior: the extractor refines the retrieved content for greater precision, while the planner enables MA-RAG to reason effectively across a range of question types.
@@
Impact of LLMs To assess the effect of model size on different agents in our multi-agent system, we conduct an ablation study where each agent is individually replaced with a smaller 8B model, while keeping the others as LLaMA3-70B. This setup allows us to isolate the impact of LLM capacity across agents on two multi-hop datasets: HotpotQA and 2WikimQA. The numerical results are shown in Table 2. We observe that replacing the QA agent consistently leads to the largest performance degradation, with a particularly noticeable drop on 2WikimQA. This suggests that high-capacity models are crucial for final answer generation, especially when reasoning across multiple documents. Substituting either the planner or extractor agent also results in clear performance declines. In the case of the extractor, this highlights the difficulty of identifying relevant evidence in complex retrieval settings and suggests that larger models are more effective at filtering pertinent information. Meanwhile, the planner’s role in generating effective reasoning plans is also sensitive to model capacity, especially in multi-hop scenarios. In contrast, reducing the size of the step definer agent leads to only marginal performance changes, indicating that its structured role may be less reliant on high-capacity models.
@@
Dataset PubmedQA MedMCQA
@@
MA-RAG 70B Plan: ["Identify Edward De Vere, 17th Earl of Oxford’s father", "Determine the name of Edward De Vere, 17th Earl of Oxford’s paternal grandfather based on the father’s name."]
@@
Extracted Information: doc-129772: [No related information about his father.], doc-6127861: [No related information about his father.], doc-129773: [Edward De Vere, 17th Earl of Oxford’s father is John de Vere, 16th Earl of Oxford. This is evidenced by the passage stating, "De Vere was the only son of John de Vere, 16th Earl of Oxford."], doc-14010350: [No related information about his father.], doc-12540430: [No related information about his father.]
@@
Limitations MA-RAG’s multi-agent design, while improving reasoning and interpretability, introduces additional runtime and token overhead. Each agent invocation involves separate prompts and responses, which can increase latency and inference cost, especially for complex queries requiring multiple reasoning steps. Although agents are called on demand to reduce unnecessary computation, the overall workflow remains more resource-intensive than single-pass or standalone RAG systems. Our ablation study shows that not all agents require large models. Exploring how to assign resource-efficient agents to different subtasks is a promising direction for future research.

### v2

@@ Abstract
We present MA-RAG, a Multi-Agent framework for Retrieval-Augmented Generation (RAG) that addresses the inherent ambiguities and reasoning challenges in complex information-seeking tasks. Unlike conventional RAG methods that rely on end-to-end fine-tuning or isolated component enhancements, MA-RAG orchestrates a collaborative set of specialized AI agents: Planner, Step Definer, Extractor, and QA Agents, each responsible for a distinct stage of the RAG pipeline. By decomposing tasks into subtasks such as query disambiguation, evidence extraction, and answer synthesis, and enabling agents to communicate intermediate reasoning via chain-of-thought prompting, MA-RAG progressively refines retrieval and synthesis while maintaining modular interpretability. Extensive experiments on multi-hop and ambiguous QA benchmarks, including NQ, HotpotQA, 2WikimQA, and TriviaQA, demonstrate that MA-RAG significantly outperforms standalone LLMs and existing RAG methods across all model scales. Notably, even a small LLaMA3-8B model equipped with MA-RAG surpasses larger standalone LLMs, while larger variants (LLaMA3-70B and GPT-4o-mini) set new state-of-the-art results on challenging multi-hop datasets. Ablation studies reveal that both the planner and extractor agents are critical for multi-hop reasoning, and that high-capacity models are especially important for the QA agent to synthesize answers effectively. Beyond general-domain QA, MA-RAG generalizes to specialized domains such as medical QA, achieving competitive performance against domain-specific models without any domain-specific fine-tuning. Our results highlight the effectiveness of collaborative, modular reasoning in retrieval-augmented systems: MA-RAG not only improves answer accuracy and robustness but also provides interpretable intermediate reasoning steps, establishing a new paradigm for efficient and reliable multi-agent RAG [^1] Our code is available at https://github.com/thangylvp/MA-RAG.
@@ 1 Introduction
Recent advances in natural language processing have driven the development of Retrieval-Augmented Generation (RAG) models, which aim to enhance the factual accuracy and contextual relevance of generated text by integrating external knowledge sources (Lewis et al., 2020; Guu et al., 2020; Izacard & Grave, 2021; Lin et al., 2024). These systems address core limitations of Large Language Models (LLMs), such as outdated knowledge (Zhang et al., 2023b; Kasai et al., 2023) and poor generalization to domain-specific queries (Siriwardhana et al., 2023; Xiong et al., 2024), by retrieving top-$k$ documents from an external corpus (Formal et al., 2022; Izacard et al., 2022; Wang et al., 2022a) to ground the model’s output in relevant evidence.
Prior research in RAG has largely concentrated on optimizing three key components—retrieval, augmentation, and generation (Gao et al., 2024; Fan et al., 2024) (Figure 1(a)). Retrieval strategies span sparse methods (Jones, 1972; Robertson & Zaragoza, 2009) and dense retrieval (Reimers & Gurevych, 2019; Karpukhin et al., 2020), each with respective weaknesses such as lexical gaps (Berger et al., 2000) or retrieval failure on out-of-distribution and multi-hop queries (Dai et al., 2023). Augmentation methods often rely on post-retrieval processing such as re-ranking or document summarization (Chen et al., 2020; Glass et al., 2022; Ma et al., 2024) (Figure 1(b)) to improve input quality for the LLM, but add latency and may still fail to filter irrelevant or misleading evidence. More sophisticated approaches introduce iterative retrieval or query rewriting (Jiang et al., 2023b; Asai et al., 2024) (Figure 1(c)), but commonly assume the input query is well-formed and overlook the broader reasoning process across the pipeline. Across retrieval, augmentation, and generation, most existing methods treat components in isolation, failing to resolve ambiguities and reasoning gaps that span multiple stages, such as vague queries, incomplete retrievals, or scattered evidence, ultimately limiting robustness and transparency in complex QA scenarios.
@@
  • Through agent-specific chain-of-thought reasoning, MA-RAG provides interpretable intermediate steps and demonstrates strong generalization to specialized domains, such as biomedical QA, without requiring domain-specific fine-tuning.
@@ 2 Related Works
Large Language Models (LLMs) have driven significant advancements in recent years. Starting with GPT-1 (Radford et al., 2018) on the Transformer architecture (Vaswani et al., 2017), subsequent models such as GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), and GPT-4 (OpenAI, 2024) have greatly enhanced capabilities in text understanding and generation. Beyond GPT, models including Mistral (Jiang et al., 2023a), Gemini (Gemini Team, 2023), and LLaMA ((Touvron et al., 2023a), Touvron et al. (2023b)) show strong performance across tasks such as question answering and entity recognition (Zhao et al., 2023). LLM training involves unsupervised pre-training, supervised fine-tuning, and alignment with human feedback, yet domain-specific challenges remain (Kandpal et al., 2023). Techniques such as PEFT (Houlsby et al., 2019a) improve fine-tuning efficiency, while prompt-based learning (Lester et al., 2021; Li & Liang, 2021), adapters (Houlsby et al., 2019b; Fu et al., 2021; Wang et al., 2022b; He et al., 2022), and reparameterization methods (Hu et al., 2022; Edalati et al., 2022; Dettmers et al., 2023) selectively adjust parameters for improved performance. Additionally, recent studies (Wei et al., 2024; Java et al., 2025) highlight the importance of evaluating factuality and retrieval efficiency in short-form and deep research scenarios.
Retrieval-Augmented Generation (RAG) enhances LLM performance by integrating external knowledge via document retrieval (Lewis et al., 2020; Guu et al., 2020). Challenges remain in determining what, when, and how to retrieve (Gao et al., 2024). Early methods incorporate retrieval into next-token prediction (Khandelwal et al., 2020; Ram et al., 2023; Liu et al., 2024b) or use end-to-end pipelines (Guu et al., 2020; Borgeaud et al., 2022; Izacard et al., 2023; Zhang et al., 2024a), while others study knowledge representation and retrieval robustness (Xu et al., 2024; Sarthi et al., 2024). Supervised and contrastive approaches often face scalability and domain transfer limitations (Dai et al., 2023; Zhang et al., 2023b; Shi et al., 2024). Structured retrieval methods include HippoRAG (Gutiérrez et al., 2024), which leverages knowledge graphs. Recent research focuses on tighter retrieval-reasoning integration. RA-DIT (Lin et al., 2024) tunes LLMs for context use and retrievers for relevance independently. Speculative RAG (Wang et al., 2025) drafts answers with a specialist LM and verifies with a generalist LM. CD-LM (Li et al., 2025b) improves inference via chunk-level retrieval, while Auto-GDA (Leemann et al., 2025) addresses domain adaptation with synthetic data. Graph-based methods such as ToG-2 (Ma et al., 2025) and SubgraphRAG (Li et al., 2025a) utilize subgraph structures to enhance retrieval. Reinforcement learning approaches optimize retrieval and generation via episodic memory (Shinn et al., 2023), policy optimization (Kulkarni et al., 2024), evidence citation (Menick et al., 2022), and reflection-based refinement (Asai et al., 2024; Zhou et al., 2023; Gao et al., 2025). Recent open-source agentic retrieval frameworks, such as Open Deep Search (Alzubi et al., 2025a), further demonstrate the potential of lightweight reasoning agents for democratized, structured search. Compared to prior agent-based RAG systems, our MA-RAG provides a training-free, efficient, and interpretable solution.
@@
LLM-based Agentic Systems coordinate multiple specialized agents to solve complex tasks through structured interaction (Guo et al., 2024). Agents operate in diverse environments, including sandboxed, physical, or abstract settings (Hong et al., 2024; Mao et al., 2025; Park et al., 2023), and assume predefined, emergent, or data-driven roles (Du et al., 2024; Xiong et al., 2023). Communication follows cooperative, competitive, or debate-based paradigms through centralized or decentralized channels (Liu et al., 2024c; Hong et al., 2024). Capabilities are developed via environmental feedback, memory retrieval, or self-evolution (Wang et al., 2023; 2024; do Nascimento et al., 2023; Zhang et al., 2023a; Chen et al., 2024a; b). Recent work has extended agentic designs to RAG. For example, Agentic RAG for time series (Ravuru et al., 2024) uses hierarchical agent routing, while CollEX (Schneider et al., 2025) enables multimodal retrieval via vision-language agents. MA-RAG differs by employing lightweight, specialized agents that collaborate through chain-of-thought reasoning, improving transparency and performance in complex QA settings without fine-tuning.
@@ 3 Method
Preliminaries Retrieval-augmented generation leverages a large corpus of documents or contexts—such as Wikipedia—to provide grounded knowledge for question answering. Given a query $q$ and a corpus $\mathcal{C}$, a dense retriever $\mathcal{R}$ retrieves the top-$k$ relevant contexts $C_{q}=\{c_{1},...,c_{k}\}$. In standard RAG pipelines, a large language model (LLM) generates the answer based on a prompt that includes the query and the retrieved documents:
@@ 3.1 Multi-Agent System for Retrieval-Augmented Generation (MA-RAG)
Planner Agent. The Planner analyzes the input query $q$ to perform query disambiguation and task decomposition. It identifies ambiguities or underspecified elements and reformulates them into clearer sub-questions if needed. For complex or multi-hop queries, it produces a structured plan $P=\{s_{1},s_{2},\ldots,s_{n}\}$, where each $s_{i}$ denotes a reasoning subtask. The number of reasoning steps is determined by the Planner, and at each step the system performs a Retrieve $\rightarrow$ Answer RAG process using a sub-query refined by the Step Definer. Breaking down questions into simpler, targeted sub-queries enables more precise retrieval, and our empirical results across benchmarks validate this design. The Planner is guided by chain-of-thought prompting with few-shot examples, ensuring interpretable step-wise decomposition that supports grounded reasoning in downstream modules.
@@
Extractor Agent. Retrieved passages often contain redundant or irrelevant content. Instead of appending entire chunks, the Extractor selects and aggregates sentences or spans directly aligned with the current subquery. This not only filters out noise and mitigates the lost-in-the-middle issue (Liu et al., 2024a), but also enables effective evidence aggregation by combining complementary information from multiple sources into a concise evidence set for the QA agent. To avoid context overflow in multi-hop queries, the Extractor summarizes relevant content at each step, and only the step-level query with the extracted summary or answer is passed forward. This preserves continuity while keeping context concise and efficient, ultimately supporting more accurate and informed answer generation.
@@
(a) Exact Match (EM) performance of MA-RAG and baseline methods on NQ, HotpotQA, and 2WikimQA. The green star indicates MA-RAG with LLaMA3-8B, the blue star indicates MA-RAG with LLaMA3-70B, and the red star indicates MA-RAG with GPT-4o-mini. Across all datasets, MA-RAG consistently outperforms baseline methods using the same model size, demonstrating the effectiveness of our multi-agent reasoning approach.
@@ 4.1 Results
Figure 5(a) provides a visual comparison between MA-RAG and several baseline models across multiple datasets. The full numerical results with additional discussion are available in the Appendix. Key observations from our experiments include:
@@
When scaling up to larger models, MA-RAG (Llama3-70B and GPT-4o-mini) outperforms the strongest 70B-scale models, such as ChatQA-1.5 70B and RankRAG 70B, setting new state-of-the-art results on multiple benchmarks. Notably, MA-RAG achieves a score of 59.5 on NQ, 87.2 on TriviaQA, 52.1 on HotpotQA, and 47.5 on 2WikimQA. In particular, on more challenging, multi-hop, and long-tailed datasets like HotpotQA and 2WikimQA, MA-RAG demonstrates significant gains over previous methods. These improvements suggest that the fine-grained query decomposition and passage extraction capabilities inherent in MA-RAG are particularly advantageous in handling complex retrieval conditions. A key strength of our modular design is that the number of steps in MA-RAG is dynamically determined by the planner based on question complexity, with multi-hop questions resulting in more steps and LLM calls. For example, on HotpotQA, MA-RAG averages 2.3 steps per question, while on NQ, which is mostly single-hop, it averages 1.4 steps. Overall, these results highlight the critical role of multi-agent coordination in improving open-domain QA performance, emphasizing that the integration of specialized agents for different reasoning steps leads to more effective and efficient utilization of external knowledge sources.
@@ 4.2 Ablation Study
Impact of Agents. To understand each agent’s contribution, we conduct an ablation study by removing either the Extractor or Planner from MA-RAG. Table 1 reports performance across five QA benchmarks. Without the Extractor, retrieved documents are fed directly into the prompt, leading to consistent performance drops and highlighting its role in refining input and grounding responses. Removing the Planner reduces MA-RAG to a single-turn RAG system with document filtering but no query decomposition. While it still performs well on simpler, single-hop datasets, it struggles with multi-hop questions requiring structured reasoning. The largest performance drop occurs on multi-hop datasets, emphasizing the Planner’s importance in guiding complex reasoning. These results show both agents are essential: the Extractor enhances precision, and the Planner enables effective reasoning across diverse question types.
@@
Impact of LLMs. To assess the effect of model size on different agents in our multi-agent system, we conduct an ablation study where each agent is individually replaced with an LLaMA3-8B model while keeping the others as LLaMA3-70B. This isolates the impact of LLM capacity across agents on two multi-hop datasets: HotpotQA and 2WikimQA (Table 2). Replacing the QA agent consistently causes the largest performance drop, especially on 2WikimQA, suggesting that high-capacity models are crucial for final answer generation. Substituting the Planner or Extractor also leads to clear declines, with the Extractor highlighting the challenge of identifying relevant evidence in complex retrieval. The Planner’s ability to generate effective reasoning plans is similarly sensitive to model capacity. In contrast, reducing the Step Definer has only marginal impact, indicating its structured role is less dependent on large models.
@@
Method PubmedQA MedMCQA
@@
MA-RAG 70B Plan: [”Identify Edward De Vere, 17th Earl of Oxford’s father”, ”Determine the name of Edward De Vere, 17th Earl of Oxford’s paternal grandfather based on the father’s name.”]
@@
Extracted Information: doc-129772: [No related information about his father.], doc-6127861: [No related information about his father.], doc-129773: [Edward De Vere, 17th Earl of Oxford’s father is John de Vere, 16th Earl of Oxford. This is evidenced by the passage stating, ”De Vere was the only son of John de Vere, 16th Earl of Oxford.”], doc-14010350: [No related information about his father.], doc-12540430: [No related information about his father.]
@@
MA-RAG with Internet Access. We further evaluated MA-RAG with web access by integrating Google Search as the retrieval engine on the SimpleQA benchmark (Wei et al., 2024). Results in the Appendix show that MA-RAG achieves strong performance (GPT-4o: 40.1% versus MA-RAG (GPT-4o-mini, web): 86.4%), demonstrating its flexibility and competitive accuracy among recent web-enabled systems.
Discussions. MA-RAG’s multi-agent design, while improving reasoning and interpretability, introduces additional runtime and token overhead. Each agent invocation involves separate prompts and responses, which can increase latency and inference cost, especially for complex queries requiring multiple reasoning steps. Although agents are called on demand to reduce unnecessary computation, the workflow remains more resource-intensive than single-pass or standalone RAG systems. On single-hop datasets, MA-RAG achieves an average response time of about 2.2 seconds using GPT-4o-mini, while on multi-hop questions the response time increases to about 4.1 seconds. We believe this latency remains acceptable for real-world use, especially given the significant performance gains achieved. Our ablation study also shows that not all agents require large models, suggesting that assigning resource-efficient agents to different subtasks is a promising direction.